A uniform bar of mass m = 0.4 kg and length L = 0.8m is articulated around one end and oscillates in a vertical plane. Find the period of oscillation if the range of motion is small. Icm = (1/12) mL ^ 3
Answer #1

We need to find moment of inertia about the pivot.

Therefore, we will use parallel axis theorem

Ipivot = Icm + md2

Ipivot = mL2 / 12 + m (L/2)2

Ipivot = mL2 / 3

for small angle ( small range of motion)

T = 2 sqrt ( I / mgh)

where h = L/2

so,

T = 2 sqrt ( 2mL2 / 3mgL)

T =   2 sqrt (2L / 3g)

T = 2 sqrt (2 * 0.8 / 3 * 9.8)

T = 1.465 sec